QUESTION 9 Assume XYZ stock is one the 30 stocks that comprise the DJIA. The closing price for XYZ on January 10, 2018 was $231.48 On January 11, 2018, the DJIA opened at 10,097.90. The divisor at that time was.1350. Throughout the course of the day, the other 29 stock prices within the index remained unchanged and XYZ increased 3.2 percent. No stock spilits or stock dividends occurred throughout the day. Find the closing value of the DJIA on September 11, 2018. Round intermediate steps and your final answer to four decimals. O 10421.0328 10152.7693 0 11867.4360 O Cannot be determined QUESTION 10 Equally weighted indexes are disproportionately influenced by stocks with high market values. O True O False QUESTION 11 Index funds and exchange-traded funds represent a cost effective way for investors to create a diversified portfolo O True O False
Answer #1

9. if XYZ increased by 3.2%, and others remains the same, the increased value will be= $231.38*0.032= $7.40416

now the stock price becomes as= $238.78416

(if we divide 7.40416 with 0.1350, then it will be= 54.84563 points will be increase)

then the value of index will be= 10097.90+54.84563= 10.152.7456

hence option B is the right answer.

10. False, equal weighted indexes reflected with same proportion changes.

11. True, these funds will provide cost effective way to investors.